# Weak–strong uniqueness for bi-fluid compressible system with algebraic closure

### v1

@@ 1 Introduction
Multi-component compressible flow models (often termed multifluid or two-phase models) arise from interface-averaging procedures applied to microscopic balance laws. In the averaged description, volume fractions become additional unknowns and various closure assumptions are needed to close the macroscopic system. Classical modeling expositions may be found in the monographs of Ishii and Hibiki [13] and Drew and Passman [7], as well as in more modern surveys of mathematical literature on multifluid systems [4, 18].
@@
  • Algebraic closure: the phases share a common velocity and are coupled by an implicit constitutive relation—most prominently equal pressure $p_{+}=p_{-}$. In this setting, $\alpha$ is not an independent transported scalar once one passes to the “transported” partial masses $R=\alpha{\varrho_{+}}$, $Q=(1-\alpha){\varrho_{-}}$; instead, the implicit pressure relation generates additional nonlinear structure, see below for more description.
The present paper analyzes the second class and considers the viscous compressible bi-fluid system in a bounded three-dimensional domain $\Omega$, where both phases are barotropic (power-law pressures) and the single momentum equation is driven by the mixture density $R+Q$.
For mono-fluid compressible Navier–Stokes equations, Lions-Feireisl theory [22, 10] relies crucially on monotonicity/compensated compactness mechanisms tied to a pressure depending on a single conserved quantity. In multifluid systems, the pressure typically depends on several quantities; even when these quantities satisfy the continuity equations, their weak convergences do not in general commute with nonlinear implicit pressure maps. This gives rise to new challenges.
@@
The fully time-dependent viscous bi-fluid model with algebraic pressure closure was then treated by Novotný and Pokorný, who proved the existence of global-in-time finite-energy weak solutions for large initial data [26]. Their strategy was to transform the original variables into a two-densities Navier–Stokes-type system with a pressure of “complicated structure dependent on two variable densities”, and then to adapt the Lions–Feireisl method supplemented by DiPerna–Lions transport/renormalization ideas[6] refined in [24, 28].
@@
These results were then extended for differential closures (transport of $\alpha$) by Novotný [25]. In particular, Jin and Novotný [15] established the weak–strong uniqueness property of such solutions, emphasizing the role of convexity of the associated Helmholtz potential. See also [17] for problems with inflow-outflow boundary conditions, and [14] for the analogous stability result but for the dissipative measure-valued solutions. Still, the assumption about convexity of the free Helmholtz energy with respect to the conservative variables seems to be of paramount importance.
@@
On the strong-solution side, local-in-time well-posedness for the compressible two-fluid algebraic-closure system was obtained by Piasecki and Zatorska using the maximal-regularity techniques [27]. The existence of global-in-time strong solutions for small perturbations of constant initial conditions is due to Evje, Wang and Wen [8].
@@
Given the coexistence of global weak solutions and local strong solutions, a natural stability question is whether weak solutions are unique as long as a strong solution exists—the weak–strong uniqueness principle. For the mono-fluid compressible Navier–Stokes system, weak–strong uniqueness via the relative entropy method is nowadays classical [11], and has proven to have profound applications in the study of singular-limits problems, see for example [12, 23]. For the two-fluid system, the only known results, as far as we are aware, are aforementioned [15, 14] and a conditional result by Li and Zatorska [21]. The singular-limit problems, including for example the low Mach number limit, are therefore still mostly formal [18], however some recent progress for multi-component flows of MHD type can be found in [9].
@@
In sharp contrast to differential-closure models, the algebraic constraint ${p_{+}}={p_{-}}$ changes the evolution of the volume fraction to transport with an additional nonlinear term. This phenomenon appears already in the Stokes analysis of Bresch–Mucha–Zatorska [3] and is precisely what breaks the most direct relative-entropy estimates: extra terms arise that do not cancel as in the pure-transport case. The goal of this paper is to prove the weak–strong uniqueness principle for the viscous compressible bi-fluid system with algebraic pressure closure: any finite-energy weak solution in the sense of Novotný–Pokorný [26] coincides, on the lifespan of the strong solution, with the local strong solution constructed by Piasecki–Zatorska [27] emanating from the same initial data. Our proof follows the relative entropy strategy but requires an additional mechanism to control the defect in the volume fraction generated by the algebraic closure. More concretely, we exploit the specific implicit structure of the pressure law to derive a stability estimate for solution compatible with the relative energy inequality and then conclude by a Grönwall argument. It should be emphasized that, in sharp contrast with Jin–Novotný [15], no positive bound for $\alpha$ from below and away from $1$ is required in our analysis.
@@ 2 Formulation of the problem
In the system above, we denoted ${\bf S}={\bf S}(\nabla{\bf u})$ to be the usual Newtonian stress tensor given by
$\operatorname{div}{\bf S}(\nabla{\bf u})=\mu\Delta{\bf u}+(\mu+\lambda)\nabla\operatorname{div}{\bf u},$
where the two constants $\mu,\lambda$ are viscosity coefficients. We take $\mu>0$ and $2\mu+3\lambda\geq 0$. By $_{+}, pp_{-}$ we denote the internal barotropic pressures for each fluid with the explicit form:
@@
where $\gamma^{\pm}>1$ are given adiabatic exponents. The system is endowed with boundary condition
@@
on $(0,T)\times\partial\Omega$, where $\Omega$ is a sufficiently smooth bounded domain of $\mathbb{R}^{3}$ and $T>0$ is arbitrary large. We consider initial conditions in $\Omega$
@@ 2.1 Global weak solutions
We now recall the definition of a weak solution to the system (2.1), see Definition 2 from Novotný and Pokorný [26], and the corresponding existence Theorem 2. For notational convenience, we set $I:=(0,T)$.
Definition 1
The quadruple $(\alpha,\varrho_{-},\varrho_{+},{\bf u})$ is called a bounded energy weak solution to the problem (2.1)–(2.4) in $I\times\Omega$, if $0\leq\alpha\leq 1$, $\varrho_{\pm}\geq 0$ $a.$ in $I\times\Omega$, $\varrho_{\pm}\in L^{\infty}(I;L^{1}(\Omega))$, ${\bf u}\in L^{2}(I;W^{1,2}_{0}(\Omega;\mathbb{R}^{3}))$, $(\alpha\varrho_{+}+(1-\alpha)\varrho_{-})|{\bf u}|^{2}\in L^{\infty}(I;L^{1}(\Omega))$, ${p_{-}}({\varrho_{-}})={p_{+}}({\varrho_{+}})\in L^{1}(I\times\Omega)$, and:
@@
Under the algebraic closure for the pressure, it has been proven by Novotný and Pokorný [26] (see Section 5 therein) that the system (2.1) admits global-in-time weak solutions.
@@
Then the problem (2.1)–(2.4) admits at least one weak solution in the sense of Definition 1. Moreover, $\alpha\varrho_{+}$ belongs to the space $C_{\rm{weak}}([0,T];L^{\gamma^{+}}(\Omega))$ and $(1-\alpha)\varrho_{-}$ belongs to the space $C_{\rm{weak}}([0,T];L^{q_{\gamma^{+},\gamma^{-}}}(\Omega))$ [^1] Here, the notation is $q_{\theta_{1},\theta_{2}}=\theta_{1}$ if $\theta_{2}<\theta_{1}$ and $q_{\theta_{1},\theta_{2}}=\theta_{2}$ if $\theta_{2}\geq\theta_{1}$., the vector field $(\alpha\varrho_{+}+(1-\alpha)\varrho_{-}){\bf u}$ belongs to $C_{\rm{weak}}([0,T];L^{r}(\Omega;\mathbb{R}^{3}))$ for some $r>1$, and $p_{\pm}(\varrho_{\pm})\in L^{r}((0,T)\times\Omega)$ for some $r>1$.
@@ 2.2 Local-in-time strong solutions
$\partial_{t}((R+Q){\bf u})+\operatorname{div}((R+Q){\bf u}\otimes{\bf u})-\operatorname{div}{\bf S}({\bf u})+\nabla Z^{\gamma^{+}}={\bf 0}.$ (2.13c)
In the system above the pressure $p$ is expressed in terms of $R,Q$. In fact we have
@@
Relation (2.14) identifies $Z$ as $\varrho_{+}$} and thus, (2.16) can be used to deduce existence of $\alpha$ such that $0\leq\alpha\leq 1$ and $R=\alpha Z$, see [3] and also [26] for a similar argument.
@@
Assume that $\Omega$ is a uniform $C^{2}$ domain and $2<p<\infty,\;3<q<\infty,\;\frac{2}{p}+\frac{3}{q}<1$. Assume moreover that $R_{0},Q_{0}$ satisfy
@@
then (2.13) supplemented by the boundary condition (2.3), the initial conditions (2.4) and the relation (2.15) admits a unique solution $(R,Q,{\bf u})$ on $(0,T)$ with the estimate
@@
$\|\partial_{t}{\bf v}+{\bf v}\cdot\nabla{\bf v}\|_{L^{1}(0,T;L^{\infty}(\Omega))}+\|{\bf v}\|_{L^{1}(0,T;W^{1,\infty}(\Omega))}\leq C,$
@@ 3 Relative entropy inequality–preliminaries
In the following parts of this section, we will aim at writing an equation for evolution of the relative energy for $\beta,{\widetilde{R}},{\widetilde{Q}}$ and $\bf v$} sufficiently regular. The distinct difference with respect to the two-fluid model of Baer-Nunziato type [15] is that in our case, the equations for $\varrho_{+}$ and $\varrho_{-}$ are not just the continuity equations, but are more complicated. This follows from the fact that $\alpha$ does not satisfy a pure transport equation, see (3.4) for more details.
@@ 3.1 Derivation of the equations for $\alpha$, $\varrho_{+}$ and $\varrho_{-}$}
Equation (3.4) is of course not exactly a pure transport equation, but we should notice that because $0\leq\alpha\leq 1$, the coefficient
@@
is also bounded in $L^{\infty}((0,T)\times\Omega)$. Indeed, $\gamma(1-\alpha)+\alpha$ has a strictly positive lower bound, namely $\gamma(1-\alpha)+\alpha\geq\min\{1,\gamma\}>0$. Thus $|\omega_{\alpha}|\leq\frac{\gamma+1}{\min\{1,\gamma\}}$.
@@
From the equation (3.4), and (2.1a), (2.1b) we now deduce
@@
In order to derive the equations (3.4) and (3.5) rigorously, it suffices to follow the arguments from Section 8.1. in [24], see also Lemma 2.4 in [3]. In particular, we have the following result.
@@
Let ${\bf u}\in L^{2}(0,T;W^{1,2}_{0}(\Omega;\mathbb{R}^{3}))$, $R=\alpha{\varrho_{+}}\in L^{\infty}(0,T;L^{2}(\Omega))$, and $Q=(1-\alpha){\varrho_{-}}\in L^{\infty}(0,T;L^{2}(\Omega))$, and let $(R,Q,{\bf u})$ solve (2.13a) and (2.13b) in the sense of distributions. Then $Z$ defined by (2.15) belongs to $L^{\infty}(0,T;L^{2}(\Omega))$ and it satisfies
@@
Let assumptions of Lemma 3.1 be satisfied. Then $\varrho_{+}$} belongs to $L^{\infty}(0,T;L^{2}(\Omega))$ and it satisfies
@@
Proof. Let us notice that equation (3.8) follows directly from (3.7) after substituting $Z={\varrho_{+}}$ and $R=\alpha{\varrho_{+}}$. Equation for $\alpha$ is derived the same way thanks to the decomposition (3.6). The bound $0\leq\alpha\leq 1$ follows from the existence result, see Theorem $2.$ $\Box$

### v2

@@ 1 Introduction
Multi-component compressible flow models (often termed multifluid or two-phase models) arise from interface-averaging procedures applied to microscopic balance laws. In the averaged description, volume fractions become additional unknowns and various closure assumptions are needed to close the macroscopic system. Classical modeling expositions can be found in the monographs of Ishii and Hibiki [18] and Drew and Passman [9], as well as in more modern surveys of the mathematical literature on multifluid systems [5, 24].
@@
  • Algebraic closure: the phases share a common velocity and are coupled by an implicit constitutive relation—most prominently equal pressure $p_{+}=p_{-}$. In this setting, $\alpha$ is not an independent transported scalar once one passes to the “transported” partial masses $R=\alpha{\varrho_{+}}$, $Q=(1-\alpha){\varrho_{-}}$; instead, the implicit pressure relation generates an additional nonlinear structure; see below for more details.
The present paper analyzes the second class and considers the viscous compressible bi-fluid system in a bounded three-dimensional domain $\Omega$, where both phases are barotropic (power-law pressures) and the single momentum equation is driven by the mixture density $R+Q$. Our main contribution is the proof of the weak–strong uniqueness principle: any finite-energy weak solution in the sense of Novotný–Pokorný [35] coincides, on the lifespan of the strong solution, with the strong solution constructed by Piasecki–Zatorska [36], if both solutions emanate from the same initial data.
For mono-fluid compressible Navier–Stokes equations, Lions-Feireisl theory [31, 14] relies crucially on monotonicity/compensated compactness mechanisms tied to a pressure depending on a single conserved quantity. In multifluid systems, the pressure typically depends on several quantities; even when these quantities satisfy the continuity equations, their weak convergences generally do not commute with nonlinear implicit pressure maps. This gives rise to new challenges.
@@
The fully time-dependent viscous bi-fluid model with algebraic pressure closure was then treated by Novotný and Pokorný, who proved the existence of global-in-time finite-energy weak solutions for large initial data [35]. Their strategy was to transform the original variables into a two-densities Navier–Stokes-type system with a pressure of “complicated structure dependent on two variable densities”, and then to adapt the Lions–Feireisl method supplemented by DiPerna–Lions transport/renormalization ideas[8] refined in [33, 39]. Notably, all of these results required certain constraint allowing comparison between the two continuity equations, which has been subsequently removed by Wen [40]. Other extensions of this theory include existence results for the compressible viscous non-resistive MHD model, for which we refer to Li and Sun [26, 27].
For the sake of completeness, let us also recall that the existence of weak solutions to the inviscid two-fluid system with algebraic pressure closure was considered in [28], where the solutions were constructed using the convex integration method. We also refer to [10, 29, 12] for the existence results of weak solutions to related models in one space-dimension case.
@@
The results from [35] were further adapted to treat differential closure (transport of $\alpha$) by Novotný [34]. For this system, Jin and Novotný [20] established the weak–strong uniqueness property of energy-bounded weak solutions, emphasizing the role of convexity of the associated Helmholtz potential. We refer also to [22] for problems with inflow-outflow boundary conditions and [19] for the weak–strong stability result but for the dissipative measure-valued solutions. Still, the assumption about convexity of the free Helmholtz energy with respect to the conservative variables seemed to be of paramount importance. Removal of this constraint is to date an important open problem that we resolved in the current paper.
@@
On the strong-solution side, local-in-time well-posedness for the compressible two-fluid algebraic-closure system was obtained by Piasecki and Zatorska using the maximal-regularity techniques [36]. The existence of global-in-time strong solutions for small perturbations of constant initial conditions for the nonconservative model is due to Evje, Wang and Wen [11].
@@
Given the coexistence of global weak solutions and local strong solutions, a natural stability question is whether weak solutions are unique as long as a strong solution exists—the weak–strong uniqueness principle. For the compressible two-fluid system, the only known results, as far as we are aware, are the aforementioned [20, 19] and a conditional result by Li and Zatorska [30].
For the mono-fluid compressible Navier–Stokes system, weak–strong uniqueness via the relative entropy method is nowadays classical [15]. Beyond uniqueness, the relative entropy inequality has proven to have profound applications in the study of singular-limits problems, and stability of numerical schemes, see, for example, [16, 32, 7]. For the viscous compressible two-fluid system, the related results are much more sparse. For the two-fluid model from [39], the incompressible and/or inviscid singular limits were considered in [38, 23, 37], with the use of the relative entropy method. We also refer to the recent result of Fanelli, Kwon and Wróblewska-Kaminska [13], where the low Mach number limit for the multi-component MHD system, in case of ill-prepared data, was proven.
The relative entropy inequality developed here has similar implications for singular limits as the pioneering result of Feireisl, Jin and Novotný [15] for mono-fluids: it supplies a robust stability mechanism that upgrades formal asymptotics to rigorous convergence, provided suitable strong limit systems are available. In particular, our method gives a direct route to justify various incompressible limits (at least for well-prepared data) studied formally and numerically, for example in [2, 24, 17]. Moreover, the method can also be extended to two-fluid systems with a differential closure relation [5, 6].
@@
In sharp contrast to the differential-closure model [20], the algebraic constraint ${p_{+}}={p_{-}}$ changes the evolution of the volume fraction to transport with an additional nonlinear term. This phenomenon appears already in the Stokes analysis of Bresch–Mucha–Zatorska [4] and is precisely what breaks the most direct relative-entropy estimates: extra terms arise that do not cancel as in the pure-transport case. Our proof introduces an additional mechanism to control the defect in the volume fraction generated by the algebraic closure. More concretely, we exploit the specific implicit structure of the pressure law to derive a stability estimate for the volume fractions, which is compatible with the relative energy inequality, and then we conclude with a Grönwall argument. It should be emphasized that, in sharp contrast with Jin–Novotný [20], no positive bound for $\alpha$ from below and away from $1$ is required in our analysis.
@@ 2 Formulation of the problem
In the above system, we denoted ${\bf S}={\bf S}(\nabla{\bf u})$ to be the usual Newtonian stress tensor given by
${\bf S}(\nabla{\bf u})=\mu\Big(\nabla{\bf u}+(\nabla{\bf u})^{T}\Big)+\lambda\operatorname{div}{\bf u}\,{\mathbf{I}},$
where we assume that $\mu>0$ and $2\mu+3\lambda\geq 0$.
By $_{+}, pp_{-}$ we denote the internal barotropic pressures for each fluid with the explicit form:
@@
where $\gamma^{\pm}>1$ are given adiabatic exponents. The system is endowed with the boundary condition
@@
in $(0,T)\times\partial\Omega$, where $\Omega$ is a sufficiently smooth bounded domain of $\mathbb{R}^{3}$ and $T>0$ is arbitrarily large, however, finite. We consider the initial conditions in $\Omega$
@@ 2.1 Global weak solutions
We now recall the definition of a weak solution to system (2.1), see Definition 2 from Novotný and Pokorný [35], and the corresponding existence Theorem 2. For notational convenience, we set $I:=(0,T)$.
Definition 2.2
The quadruple $(\alpha,\varrho_{-},\varrho_{+},{\bf u})$ is called a bounded energy weak solution to the problem (2.1)–(2.4) in $I\times\Omega$, if
$0\leq\alpha\leq 1,\quad\varrho_{\pm}\geq 0\quad\text{a.a. in }\,I\times\Omega,$
$\varrho_{\pm}\in L^{\infty}(I;L^{1}(\Omega)),\quad{\bf u}\in L^{2}(I;W^{1,2}_{0}(\Omega;\mathbb{R}^{3}),$
$(\alpha\varrho_{+}+(1-\alpha)\varrho_{-})|{\bf u}|^{2}\in L^{\infty}(I;L^{1}(\Omega)),\quad{p_{-}}({\varrho_{-}})={p_{+}}({\varrho_{+}})\in L^{1}(I\times\Omega),$
and
@@
Under the algebraic closure for the pressure, Novotný and Pokorný [35] (see Section 5 therein) have shown that the system (2.1) admits global-in-time weak solutions.
@@
Then problem (2.1)–(2.4) admits at least one weak solution in the sense of Definition $2.$ Moreover, $\alpha\varrho_{+}$ belongs to the space $C_{\rm{weak}}([0,T];L^{\gamma^{+}}(\Omega))$ and $(1-\alpha)\varrho_{-}$ belongs to the space $C_{\rm{weak}}([0,T];L^{q_{\gamma^{+},\gamma^{-}}}(\Omega))$ [^1] Here, the notation is $q_{\theta_{1},\theta_{2}}=\theta_{1}$ if $\theta_{2}<\theta_{1}$ and $q_{\theta_{1},\theta_{2}}=\theta_{2}$ if $\theta_{2}\geq\theta_{1}$., the vector field $(\alpha\varrho_{+}+(1-\alpha)\varrho_{-}){\bf u}$ belongs to $C_{\rm{weak}}([0,T];L^{r}(\Omega;\mathbb{R}^{3}))$ for some $r>1$, and $p_{\pm}(\varrho_{\pm})\in L^{r}((0,T)\times\Omega)$ for some $r>1$.
@@ 2.2 Local-in-time strong solutions
$\partial_{t}((R+Q){\bf u})+\operatorname{div}((R+Q){\bf u}\otimes{\bf u})-\operatorname{div}{\bf S}(\nabla{\bf u})+\nabla Z^{\gamma^{+}}={\bf 0}.$ (2.13c)
In the above system, the pressure $p$ is expressed in terms of $R,Q$. In fact, we have
@@
Relation (2.14) identifies $Z$ as $\varrho_{+}$} and thus (2.16) can be used to deduce the existence of $\alpha$ such that $0\leq\alpha\leq 1$ and $R=\alpha Z$, see [4] and also [35] for a similar argument.
@@
Assume that $\Omega$ is a uniform $C^{2}$ domain and $2<p<\infty,\;3<q<\infty,\;\frac{2}{p}+\frac{3}{q}<1$. Assume, moreover, that $R_{0},Q_{0}$ satisfy
@@
then (2.13) supplemented by the boundary condition (2.3), the initial conditions (2.4) and the relation (2.15) admits a unique solution $(R,Q,{\bf u})$ in $(0,T)$ with the estimate
@@
$\|\partial_{t}{\bf u}+{\bf u}\cdot\nabla{\bf u}\|_{L^{2}(0,T;L^{\infty}(\Omega))}+\|{\bf u}\|_{L^{2}(0,T;W^{1,\infty}(\Omega))}\leq C,$
@@ 3 Relative entropy inequality–preliminaries
In the following parts of this section, we will aim at writing an equation for the evolution of the relative energy for $\beta,{\widetilde{R}},{\widetilde{Q}}$ and $\bf v$} sufficiently regular. The distinct difference with respect to the two-fluid model of Baer-Nunziato type [20] is that in our case the equations for $\varrho_{+}$ and $\varrho_{-}$ are not just continuity equations, but are more complicated. This follows from the fact that $\alpha$ does not satisfy a pure transport equation, see (3.4) for more details.
@@ 3.1 Derivation of the equations for $\alpha$, $\varrho_{+}$ and $\varrho_{-}$}
Equation (3.4) is, of course, not exactly a pure transport equation, but we should notice that because $0\leq\alpha\leq 1$, the coefficient
@@
is also bounded in $L^{\infty}((0,T)\times\Omega)$. Indeed, $\gamma(1-\alpha)+\alpha$ has a strictly positive lower bound, namely $\gamma(1-\alpha)+\alpha\geq\min\{1,\gamma\}>0$. Thus $|\omega_{\alpha}|\leq\frac{|\gamma-1|}{4\min\{1,\gamma\}}<\frac{\gamma+1}{\min\{1,\gamma\}}$.
@@
From equations (3.4), and (2.1a), (2.1b), we now deduce
@@
In order to derive equations (3.4) and (3.5) rigorously, it suffices to follow the arguments from Section 8.1. in [33], see also Lemma 2.4 in [4]. In particular, we have the following result.
@@
Let ${\bf u}\in L^{2}(0,T;W^{1,2}_{0}(\Omega;\mathbb{R}^{3}))$, $R=\alpha{\varrho_{+}}\in L^{\infty}(0,T;L^{2}(\Omega))$, and $Q=(1-\alpha){\varrho_{-}}\in L^{\infty}(0,T;L^{2}(\Omega))$, and let $(R,Q,{\bf u})$ solve (2.13a) and (2.13b) in the sense of distributions. Then $Z$ defined by (2.15) belongs to $L^{\infty}(0,T;L^{2}(\Omega))$ and satisfies
@@
Let the assumptions of Lemma 3.1 be satisfied. Then $\varrho_{+}$} belongs to $L^{\infty}(0,T;L^{2}(\Omega))$ and it satisfies
@@
Proof. Let us note that equation (3.8) follows directly from (3.7) after substituting $Z={\varrho_{+}}$ and $R=\alpha{\varrho_{+}}$. The equation for $\alpha$ is derived in the same way thanks to the decomposition (3.6). The bound $0\leq\alpha\leq 1$ follows from the existence result, see Theorem 2.3. $\Box$